# GENIE: Watermarking Graph Neural Networks for Link Prediction

### v1

@@ 4 GENIE
LP using DNNs is a complex task that can be approached from various angles, which makes watermarking techniques less straightforward to implement as compared to other domains. For instance, LP task can be converted into a binary graph classification task by creating subgraphs for each link, passing each subgraph through a GNN for feature propagation, and then learning a graph-level readout function to predict a link’s existence. On the other hand, we can also train a GNN without explicit subgraph construction, i.e., by learning node representations and passing the representations through a Multi-layer Perceptron (MLP) to predict a link’s existence. The possibility of implementing LP task via different methodologies makes it difficult to devise a unified scheme to watermark LP models based on GNNs. To address this challenge, we propose GENIE to watermark GNNs for LP task. GENIE focuses on two key methods of LP in GNNs. We first focus on GNN-based models that operate directly on the graph structure; here we modify the node representations to embed the watermark. We next target the subgraph-based approaches; here we embed the watermark within the subgraph representations learned by the GNN. We describe the process of generating and embedding watermark using GENIE in §4.1 and §4.2, respectively. §4.3 and §4.4 explain the process of watermark verification and ownership demonstration, respectively, in GENIE.
@@ 4.1 Watermark data generation
Given a pair of nodes, $\mathcal{G}$, and feature vectors of all the nodes in $\mathcal{G}$, let $\mathcal{F}$ be the ground truth function that correctly classifies the existence of a link in $\mathcal{G}$. We define $\mathcal{F}_{wm}$ as a function that behaves the same as $\mathcal{F}$ for “normal” inputs but, outputs the opposite of $\mathcal{F}$ on “backdoored” or “watermark” inputs. Formally, $\mathcal{F}_{wm}$ is the ground truth function that correctly classifies the existence of a link in $\mathcal{G}_{wm}$ (cf. §4.1.1). We call $\mathcal{F}_{wm}$ as the watermark function. Our task is to make a model learn $\mathcal{F}_{wm}$ using a modified $\mathcal{D}_{train}$ and $\mathcal{D}_{wm}$ that contains the “watermark” inputs. Let $\mathcal{M}_{clean}$ be a model trained to learn $\mathcal{F}$ using $\mathcal{D}_{train}$. In order to learn the function $\mathcal{F}_{wm}$, $\mathcal{D}_{wm}$ needs to be defined in such a way that $\forall~{}(\mathcal{D}_{train},~{}\mathcal{M}_{clean})$
@@
We denote the Area Under the ROC Curve (AUC) score of $\mathcal{M}$ evaluated $\mathcal{D}$ as $AUC_{\mathcal{D}}^{\mathcal{M}}$. Figure 1 depicts the behaviour of $\mathcal{M}_{wm}$ on normal and watermark inputs.
Figure 1: The predictions by a watermarked GNN on an input graph injected with watermark (i.e., Link exists prediction) should be opposite of that on an input graph without watermark injection (i.e., Link doesn’t exist prediction).
@@ 4.1.1 GENIE for node representation-based method
Let $\mathbf{E}_{wm}\in{\mathbb{N}}^{{|\mathcal{E}_{\mathcal{S}}\cup{\overline{\mathcal{E}}}_{\mathcal{S}}|}\times 2}$ be the edge index matrix containing both the links present in ¯${\overline{\mathcal{E}}}_{\mathcal{S}}$ (labeled as positive) and the links present in $\mathcal{E}_{\mathcal{S}}$ (labeled as negative) for some arbitrary ordering of ${\mathcal{E}}_{\mathcal{S}}\cup{\overline{\mathcal{E}}}_{\mathcal{S}}$. Let $\mathbf{y}_{wm}$ be the label vector corresponding to the links in $\mathbf{E}_{wm}$. Training the model with modified $\mathbf{A}_{wm}$ and $\mathbf{X}$ directly might result in convergence difficulties and suboptimal performance. It is so because the model would get confused by the contradictory information introduced during the watermarking process (i.e., previously existing links are now absent and previously non-existing links are now present). At the same time, polarities of gradients for these modified links would be reversed, making the loss function harder to optimize. It may potentially cause the model to diverge or converge to a suboptimal solution. To address this issue, we modify the node feature matrix $\mathbf{X}$ along with $\mathbf{A}$ as follows to provide more information about the watermark.
For all nodes $v\in\mathcal{S}$, we replace the original node feature vectors $\mathbf{x}_{v}$ with a watermark vector $\mathbf{w}\in\mathbb{R}^{1\times d}$, i.e., $\mathbf{x}_{v}=\mathbf{w}$. We chose the elements of $\mathbf{w}$ from a uniform distribution. We denote the modified node feature matrix as $\mathbf{X}_{wm}$. Our intuition behind this approach is that a GNN will learn to associate the presence of watermark vector $\mathbf{w}$ in the node features with a specific LP behavior. In particular, whenever the node features of two nodes involved in a LP task are equal to $\mathbf{w}$, GNN should predict the opposite of the link’s true existence, effectively embedding the watermark information in the model’s predictions. We define the watermarking dataset as a 4-tuple $\mathcal{D}_{wm}=(\mathbf{E}_{wm},\mathbf{A}_{wm},\mathbf{X}_{wm},\mathbf{y}_{wm})$, which is used in the watermark embedding process. Similarly, training dataset is defined as $\mathcal{D}_{train}=(\mathbf{E}_{train},\mathbf{A},\mathbf{X},\mathbf{y})$, where $\mathbf{E}_{train}$ is the edge index matrix containing the links present in $\mathcal{E}_{train}$ (cf. §2.1) and $\mathbf{y}$ is the label vector corresponding to $\mathbf{E}_{train}$. We provide the details of the watermark embedding process in § 4.2.
@@ 4.1.2 GENIE for subgraph-based method
In this method, the GNN takes a subgraph as input and performs binary graph classification for LP. $\mathcal{D}_{train}$ is created by constructing a subgraph of $k$-hops around each positive link and an equal number of sampled negative links in the original graph $\mathcal{G}$, and assigning a label of 00 or $1$ (based on the existence of the link). Let $T$ denote all the subgraphs present in $\mathcal{D}_{train}$. Therefore, $\mathcal{D}_{train}=(\mathbf{G}_{T},\mathbf{y}_{clean})$, where $\mathbf{G}_{T}$ denotes the tensor collecting all the subgraphs and $\mathbf{y}_{clean}\in\{0,~{}1\}^{T}$ is the label vector for all the corresponding subgraphs. To generate $\mathcal{D}_{wm}$, we selectively modify a sample of $k$-hop subgraphs constructed from $\mathcal{G}$ (instead of modifying the entire graph $\mathcal{G}$). In particular, we first sample $s=\left\lceil\alpha T\right\rceil$ subgraphs $\mathcal{G}_{i},i\in[s]$ uniformly at random from subgraphs present in $\mathbf{G}_{T}$ and flip labels of the sampled subgraphs. Here, $\alpha\in(0,~{}1)$ is the watermarking rate. Formally, if ${\mathbf{y}}_{S}\in\{0,~{}1\}^{s}$ denotes the vector of labels of subgraphs $\mathcal{G}_{i},i\in[s]$, then $\mathbf{y}_{wm}={\overline{\mathbf{y}}}_{S}$. Next, we construct $\mathcal{D}_{wm}=(\mathbf{G}_{S},\mathbf{y}_{wm})$, where ${\mathbf{G}_{S}}$ denotes the tensor collecting the modified subgraphs $\mathcal{G}_{i},i\in[s]$. Similar to the previous method, only inverting the labels would confuse the model. To solve this issue, we replace the node feature of each node in all the subgraphs in $\mathcal{D}_{wm}$ with the watermark vector $\mathbf{w}$ (defined in §4.1.1). Figure 3 illustrates the modifications made to $\mathcal{G}_{i}$ to obtain $\mathcal{D}_{wm}$. The intuition remains the same here as well, i.e., GNN will be able to associate the presence of $\mathbf{w}$ with the inversion of labels.
Figure 3: A representative illustration of generating $\mathcal{D}_{wm}$ for subgraph-based methods. Here, the original subgrapph $\mathcal{G}_{i}$ is created from an arbitrary pair of nodes $(u,v)$ with label $y_{i}\in\{0,~{}1\}$. In the modified subgraph, the original feature vectors $\mathbf{x_{v}}$ are replaced with the watermark vector $\mathbf{w}$ and the subgraph label $\overline{y}_{i}$.
@@ 4.2 Watermark embedding
[…]
Update Step for $\mathcal{L}_{train}$:
[…]
Update Step for $\mathcal{L}_{wm}$:
[…]
Combined Update Step for each iteration $t$:
[…]
Our approach of backpropagating losses from $\mathcal{D}_{train}$ and $\mathcal{D}_{wm}$ separately is motivated by the intuition that it allows the model to effectively learn both distributions independently. By first updating $\theta$ based on $\mathcal{D}_{train}$ allows the model to capture the inherent patterns and relationships present in the non-watermarked instances. Then, updating $\theta$ based on $\mathcal{D}_{wm}$ enables the model to associate the watermark patterns with the predetermined incorrect predictions. Consequently, it embeds the watermark information effectively. Further nuances in §6.
@@ 4.3 Watermark verification
Following the works [19, 47, 38], we employ a statistical-cum-empirical approach in lieu of a theoretical approach to give guarantees of GENIE satisfying the watermarking requirements (cf. §3). In particular, we expect AUC score to be high when we query a $\mathcal{M}_{wm}$ with the secret $\mathcal{D}_{wm}$. This premise assumes $\mathcal{M}_{wm}$ to have statistically different behavior than $\mathcal{M}_{clean}$ when queried with $\mathcal{D}_{wm}$. In other words, ensuring that there is always a statistically significant difference between $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ is a must for the watermark to work.
To provide such a statistical guarantee, we use the Smoothed Bootstrap Approach (SBA) [48]. The reason to choose this test instead of conventional hypothesis tests (i.e., the parametric Welch’s t-test [49] or the non-parametric Mann-Whiteney U test [50]) is twofold. First, SBA generalizes to non-normal data. Thus, it can be used in situations where all AUC scores are not normally distributed. SBA is applicable to our data as the Shapiro-Wilk Test [51] with a significance level of $\alpha=0.05$ for some of our scores rejects the null hypothesis of the data being normally distributed [^5] The authors in [19, 47] consider data to be normally distributed.. Second, the Mann-Whiteney U test (among other non-parametric hypothesis tests) is a test to establish stochastic inequality of the distribution of the given two samples. In our case, it is trivially seen as all $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ are found to be greater than $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$. It means that performing non-parametric tests would give a positive result supporting our claim, i.e., the trivial $p$-value of 00, in all cases.
Let $\mathcal{W}_{i}$ and $\mathcal{C}_{i}$, $i\in[n]$ denote $n$ watermarked and clean models, respectively. Let $\alpha_{i}$ and $\beta_{i}$, $i\in[n]$ denote the AUC scores of models $\mathcal{W}_{i}$ and $\mathcal{C}_{i}$ on $\mathcal{D}_{wm}$, respectively. Our goal is to provide a statistical guarantee that the difference between $\overline{\alpha}$ and $\overline{\beta}$ is significant, where $\overline{\alpha}$ and $\overline{\beta}$ denote the mean of $\alpha_{i}$ and $\beta_{i}$, $i\in[n]$, respectively. To this end, our null hypothesis $\mathbf{H_{0}}$ and alternate hypothesis $\mathbf{H_{a}}$ are shown in Eq. 9 and Eq. 10, respectively:
@@
The theoretical analysis above yields a condition under which $\mathcal{O}$ can reject $\mathbf{H_{0}}$ with $\tau$ confidence level (or, $1-\tau$ significance level). We verify our analysis by assessing SEAL architecture’s performance on Yeast dataset [52]; TABLE II present the results for $n=10$. Applying the Shapiro-Wilk test on these results, we get the $p$-values of $\beta_{i}$ and $\alpha_{i}$, $i\in[n]$ to be $0.001$ and $0.339$, respectively. With a significance level of $0.05$, we reject the null hypothesis of $\beta_{i}$, $i\in[10]$ to be normally distributed. We perform smoothed bootstrap with number of bootstrap samples equal to $10^{5}$, bandwidth set according to Silverman’s rule of thumb [53] and get the $p$-value to be $0.000$. With the significance level of $0.05$, we reject the hypothesis $\mathbf{H_{0}}$ for SEAL architecture on the Yeast dataset. The results for every model architecture and dataset considered in our work are presented in Appendix B; where for each architecture-dataset combination, we reject $\mathbf{H_{0}}$ with the significance level of $0.05$ as well. Thus, GENIE satisfies the non-trivial ownership requirement (described in §3).
Using the statistical analysis similar to above, we set the thresholds for each dataset and model architecture listed in TABLE III. For the sake of brevity, we elucidate the procedure of setting the thresholds in Appendix C.

### v3

@@ 4 Genie
As evident in §2.2, the diverse GNN-based methods (e.g., SEAL, GCN) for LP present a challenge for developing a generic watermarking technique for LP. Designing a specific watermarking scheme for each variant would be impractical. However, Genie circumvents this issue by focusing on the construction of a custom dataset (i.e., $\mathcal{D}_{wm}$) which depends solely on the input structure to the GNN, rather than the specific model architecture. Genie offers a watermarking framework for both node-representation and subgraph-based approaches. We describe the process of generating and embedding watermark using Genie in §4.1 and §4.2, respectively. §4.3 and §4.4 explain the process of Genie’s watermark verification and OD, respectively.
@@ 4.1 Watermark data generation
Given a pair of nodes, $\mathcal{G}$, and feature vectors of all the nodes in $\mathcal{G}$, let $\mathcal{F}$ be the ground truth function that correctly classifies the existence of a link in $\mathcal{G}$. We define $\mathcal{F}_{wm}$ as a function that behaves the same as $\mathcal{F}$ for “normal” inputs but, outputs the opposite of $\mathcal{F}$ on “backdoored” or “watermark” inputs. Formally, $\mathcal{F}_{wm}$ is the ground truth function that correctly classifies the existence of a link in $\mathcal{G}_{wm}$ (cf. §4.1.1). We call $\mathcal{F}_{wm}$ as the watermark function. Our task is to make a model learn $\mathcal{F}_{wm}$ using a modified $\mathcal{D}_{train}$ and $\mathcal{D}_{wm}$ that contains the “watermark” inputs. Let $\mathcal{M}_{clean}$ be a model trained to learn $\mathcal{F}$ using $\mathcal{D}_{train}$. In order to learn the function $\mathcal{F}_{wm}$, $\mathcal{D}_{wm}$ needs to be created in such a way that $\forall~{}(\mathcal{D}_{train},~{}\mathcal{M}_{clean})$
@@
We denote the Area Under the ROC Curve (AUC) score of $\mathcal{M}$ evaluated against $\mathcal{D}$ as $AUC_{\mathcal{D}}^{\mathcal{M}}$. Figure 1 depicts the behaviour of $\mathcal{M}_{wm}$ on normal and watermark inputs.
Figure 1: The predictions by a watermarked GNN on a graph without watermark (i.e., Link doesn’t exist) should be opposite to that on a graph injected with watermark (i.e., Link exists).
@@ 4.1.1 Genie for node representation-based method
Let $\mathbf{E}_{wm}\in{\mathbb{N}}^{{|\mathcal{E}_{\mathcal{S}}\cup{\overline{\mathcal{E}}}_{\mathcal{S}}|}\times 2}$ be the edge index matrix containing both the links present in ¯${\overline{\mathcal{E}}}_{\mathcal{S}}$ (labeled as positive) and the links present in $\mathcal{E}_{\mathcal{S}}$ (labeled as negative) for some arbitrary ordering of ${\mathcal{E}}_{\mathcal{S}}\cup{\overline{\mathcal{E}}}_{\mathcal{S}}$. Let $\mathbf{y}_{wm}$ be the label vector corresponding to the links in $\mathbf{E}_{wm}$. Training the model with modified $\mathbf{A}_{wm}$ and $\mathbf{X}$ directly might result in convergence difficulties and suboptimal performance. It is so because the model would get confused by the contradictory information introduced during the watermarking process (i.e., previously existing links are now absent, and previously non-existing links are now present). At the same time, the polarities of gradients for these modified links would be reversed, making the loss function harder to optimize. It may potentially cause the model to diverge or converge to a suboptimal solution. To address this issue, we modify the node feature matrix $\mathbf{X}$ along with $\mathbf{A}$ as follows to provide more information about the watermark.
For all nodes $v\in\mathcal{S}$, we replace the original node feature vectors $\mathbf{x}_{v}$ with a watermark vector $\mathbf{w}\in\mathbb{R}^{1\times d}$, i.e., $\mathbf{x}_{v}=\mathbf{w}$. We chose the elements of $\mathbf{w}$ from a uniform distribution. We denote the modified node feature matrix as $\mathbf{X}_{wm}$. Our intuition behind this approach is that a GNN will learn to associate the presence of watermark vector $\mathbf{w}$ in the node features with a specific LP behavior. In particular, whenever the node features of two nodes involved in an LP task are equal to $\mathbf{w}$, GNN should predict the opposite of the link’s true existence, effectively embedding the watermark information in the model’s predictions. We define the watermarking dataset as a 4-tuple $\mathcal{D}_{wm}=(\mathbf{E}_{wm},\mathbf{A}_{wm},\mathbf{X}_{wm},\mathbf{y}_{wm})$, which is used in the watermark embedding process. Similarly, training dataset is defined as $\mathcal{D}_{train}=(\mathbf{E}_{train},\mathbf{A},\mathbf{X},\mathbf{y})$, where $\mathbf{E}_{train}$ is the edge index matrix containing the links present in $\mathcal{E}_{train}$ (cf. §2.1) and $\mathbf{y}$ is the label vector corresponding to $\mathbf{E}_{train}$. We provide the details of the watermark embedding process in § 4.2.
@@ 4.1.2 Genie for subgraph-based method
In this method, the GNN takes a subgraph as input and performs binary graph classification for LP. $\mathcal{D}_{train}$ is created by constructing a subgraph of $k$-hops around each positive link and an equal number of sampled negative links in the original graph $\mathcal{G}=(\mathcal{V},~{}\mathcal{E})$, and assigning a label of 00 or $1$ (based on the existence of the link). Let $T$ denote all the subgraphs present in $\mathcal{D}_{train}$. Therefore, $\mathcal{D}_{train}=(\mathbf{G}_{T},\mathbf{y}_{clean})$, where $\mathbf{G}_{T}$ denotes the tensor collecting all the subgraphs and $\mathbf{y}_{clean}\in\{0,~{}1\}^{T}$ is the label vector for all the corresponding subgraphs. To generate $\mathcal{D}_{wm}$, we selectively modify a sample of $k$-hop subgraphs constructed from $\mathcal{G}$ (instead of modifying the entire graph $\mathcal{G}$). In particular, we first sample $s=\left\lceil\alpha_{sg}T\right\rceil$ subgraphs $\mathcal{G}_{i},i\in[s]$ uniformly at random from subgraphs present in $\mathbf{G}_{T}$ and invert labels of the sampled subgraphs. Here, $\alpha_{sg}\in(0,~{}1)$ is the watermarking rate for subgraph-based method. Formally, if ${\mathbf{y}}_{S}\in\{0,~{}1\}^{s}$ denotes the vector of labels of subgraphs $\mathcal{G}_{i},i\in[s]$, then $\mathbf{y}_{wm}={\overline{\mathbf{y}}}_{S}$. Next, we construct $\mathcal{D}_{wm}=(\mathbf{G}_{S},\mathbf{y}_{wm})$, where ${\mathbf{G}_{S}}$ denotes the tensor collecting the modified subgraphs $\mathcal{G}_{i},i\in[s]$. Similar to the previous method, only inverting the labels would confuse the model. To solve this issue, we replace the node feature of each node in all the subgraphs in $\mathcal{D}_{wm}$ with the watermark vector $\mathbf{w}$ (defined in §4.1.1). Figure 3 illustrates the modifications made to $\mathcal{G}_{i}$ to obtain $\mathcal{D}_{wm}$. The intuition remains the same here as well, i.e., GNN will be able to associate the presence of $\mathbf{w}$ with the inversion of labels.
Figure 3: A depiction of generating $\mathcal{D}_{wm}$ for subgraph-based methods. Here, the original subgrapph $\mathcal{G}_{i}$ is created from an arbitrary pair of nodes $(u,v)$ with label $y_{i}\in\{0,~{}1\}$. In the modified subgraph, the original feature vectors $\mathbf{x_{v}}$ are replaced with the watermark vector $\mathbf{w}$ and the subgraph label $\overline{y}_{i}$.
@@ 4.2 Watermark embedding
Watermark embedding is a crucial step in any watermarking scheme, which directly affects the functionality-preserving property. Previous works have proposed various watermark embedding methods (e.g., data poisoning, multi-task learning). However, our experiments (cf. §5.3) show that the existing watermark embedding methods are sub-optimal. To this end, we introduce a novel watermark embedding method that outperforms or competes with SotA methods.
The watermark embedding process in Genie is similar for both node representation-based and subgraph-based LP methods. Since $\mathcal{D}_{train}$ is different in both methods, the nature of $\mathcal{D}_{wm}$ is also different in each of them. In particular, $\mathcal{D}_{train}$ contains links with positive or negative labels (depending on a given link’s existence) for node-representation based method. On another side, $\mathcal{D}_{train}$ contains $k$-hop subgraphs constructed around a link with positive or negative labels (depending on a given link’s existence) for subgraph-based method. To embed the watermark, we initially take an untrained GNN model $\mathcal{M}$ and train it using a combination of $\mathcal{D}_{train}$ and $\mathcal{D}_{wm}$. The training is done in a specific manner to ensure that the model effectively learns distributions from both $\mathcal{D}_{train}$ and $\mathcal{D}_{wm}$. More formally, let $\mathcal{L}_{train}$ and $\mathcal{L}_{wm}$ be the loss functions corresponding to $\mathcal{D}_{train}$ and $\mathcal{D}_{wm}$, respectively. We start with initial parameters $\bm{\theta}_{\textbf{0}}$ and learning rate $\eta$. In each training epoch, we update the model parameters as follows:
  1. 1.
    Compute the gradient of $\mathcal{L}_{train}$ with respect to the model parameters ($\bm{\theta}$) using $\mathcal{D}_{train}$.
  2. 2.
    Update $\bm{\theta}$ by applying the gradients from step (1) using an optimizer.
  3. 3.
    Compute gradient of $\mathcal{L}_{wm}$ with respect to $\bm{\theta}$ using $\mathcal{D}_{wm}$.
  4. 4.
    Update $\bm{\theta}$ by applying the gradients from step (3) using the same optimizer.
Our approach of backpropagating losses from $\mathcal{D}_{train}$ and $\mathcal{D}_{wm}$ separately is motivated by the intuition that it allows the model to effectively learn both distributions independently. By first updating $\bm{\theta}$ based on $\mathcal{D}_{train}$ allows the model to capture the inherent patterns and relationships present in the non-watermarked instances. Then, updating $\theta$ based on $\mathcal{D}_{wm}$ enables the model to associate the watermark patterns with the predetermined incorrect predictions. Consequently, it embeds the watermark information effectively.
@@ 4.3 Watermark verification
The watermark verification process requires testing $\mathcal{D}_{wm}$ against $\mathcal{M}_{adv}$. We expect Genie to have the property of non-trivial ownership, i.e., if $\mathcal{M}_{adv}$ is $\mathcal{M}_{wm}$, then $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ should be high, otherwise $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ should be low. Therefore, the watermark verification process requires: (1) an assurance of Genie satisfying the property of non-trivial ownership; and (2) a procedure to calculate the specific threshold as to how high or low should $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ be, to classify $\mathcal{M}_{adv}$ as $\mathcal{M}_{wm}$.
We first demonstrate the non-trivial ownership property of Genie using statistical hypothesis testing in §4.3.1. We then outline problems with existing watermark threshold setting procedure as a motivation for DWT (our novel threshold setting procedure) and then describe its mechanism in §4.3.2.
4.3.1 Non-trivial ownership
Following the works [12, 36, 27], we employ a statistical-cum-empirical approach in lieu of a theoretical approach to give assurance of Genie satisfying the watermarking requirements (cf. §3). To provide such a statistical assurance, we use the Smoothed Bootstrap Approach (SBA) [37]. The reason to choose this test instead of conventional hypothesis tests (i.e., the parametric Welch’s t-test [38] or the non-parametric Mann-Whiteney U test [39]) is twofold. First, SBA generalizes to non-normal data. Thus, it can be used in situations where all AUC scores are not normally distributed. SBA is applicable to our data as the Shapiro-Wilk Test [40] with a significance level of $\alpha=0.05$ for some of our scores rejects the null hypothesis of the data being normally distributed [^4] The authors in [12, 36] consider data to be normally distributed.. Second, the Mann-Whiteney U test (among other non-parametric hypothesis tests) is a test to establish stochastic inequality of the distribution of the given two samples. In our case, it is trivially seen as all $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ are found to be greater than $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$. It means that performing non-parametric tests would give a positive result supporting our claim, i.e., the trivial $p$-value of 00, in all cases.
Let $\mathcal{W}_{i}$ and $\mathcal{C}_{i}$, $i\in[n]$ denote $n$ watermarked and clean models, respectively. Let $\alpha_{i}$ and $\beta_{i}$, $i\in[n]$ denote the AUC scores of models $\mathcal{W}_{i}$ and $\mathcal{C}_{i}$ on $\mathcal{D}_{wm}$, respectively. Our goal is to provide a statistical assurance that the difference between $\overline{\alpha}$ and $\overline{\beta}$ is significant, where $\overline{\alpha}$ and $\overline{\beta}$ denote the mean of $\alpha_{i}$ and $\beta_{i}$, $i\in[n]$, respectively. To this end, our null hypothesis $\mathbf{H_{0}}$ and alternate hypothesis $\mathbf{H_{a}}$ are shown in Eq. 4 and Eq. 5, respectively:
@@ 4.3.1 Non-trivial ownership
The theoretical analysis above yields a condition under which $\mathcal{O}$ can reject $\mathbf{H_{0}}$ with $\tau$ confidence level (or, $1-\tau$ significance level). We verify our analysis by assessing SEAL architecture’s performance on Yeast dataset [41]; TABLE II present the results for $n=10$. Applying the Shapiro-Wilk test on these results, we get the $p$-values of $\beta_{i}$ and $\alpha_{i}$, $i\in[n]$ to be $0.001$ and $0.339$, respectively. With a significance level of $0.05$, we reject the null hypothesis of $\beta_{i}$, $i\in[10]$ to be normally distributed. We perform smoothed bootstrap with the number of bootstrap samples equal to $10^{5}$, bandwidth set according to Silverman’s rule of thumb [42], and get the $p$-value to be $0.000$. With the significance level of $0.05$, we reject the hypothesis $\mathbf{H_{0}}$ for SEAL architecture on the Yeast dataset. The results for every model architecture and dataset considered in our work are presented in Appendix C; where for each architecture-dataset combination, we reject $\mathbf{H_{0}}$ with the significance level of $0.05$ as well. Thus, Genie satisfies the non-trivial ownership requirement (described in §3).
4.3.2 Dynamic Watermark Thresholding (DWT)
Existing watermark threshold setting procedures can mainly be classified into 3 types [8]: (1) choosing the highest $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ as the threshold; (2) choosing the lowest $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ as the threshold; (3) averaging $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ of multiple $\mathcal{M}_{clean}$ and $\mathcal{M}_{wm}$ models. Procedures (1) and (2) will have high FPR and FNR, respectively. Procedure (3) tries to balance the trade-off of FPR and FNR present in procedures (1) and (2), but it does not give any statistical assurance. Regardless, all these procedures suffer from various other problems, like the lack of efficiency, requiring to train as many as 400 independent models (e.g., in work [7]), lack of theoretical or statistical assurance (e.g., in works [8, 7]), lack of generalizability to other watermarking schemes (e.g., in work [9]), and the naive assumption of normality of data (e.g., in works [12, 36, 10]). These problems greatly limit a watermarking scheme’s practical applicability. Considering these issues, we identify the following four main properties of an ideal watermark threshold setting procedure:
  1. 1.
    Efficiency: The watermark threshold setting procedure should utilize as few $\mathcal{M}_{clean}$ and $\mathcal{M}_{wm}$ as possible.
  2. 2.
    Assurance: The procedure should be able to give assurance of its correctness, either theoretically or statistically.
  3. 3.
    Generality: The procedure should: (a) generalize to all kinds of watermarking scheme and model architectures; and (b) make as few assumptions about the data as possible, i.e., it should generalize to all distributions of data (viz., normal, non-normal).
  4. 4.
    Robustness: The watermark thresholds obtained from the procedure should be robust to outliers present in the data, i.e., in $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$.
To the best of our knowledge, none of the existing works have a watermark threshold setting procedure that complies to all of the aforementioned properties. We introduce DWT - a simple, yet effective watermark threshold setting procedure that encapsulates all the four identified properties.
The general idea of DWT is to use Kernel Density Estimate (KDE) to estimate the distribution of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$. Informally, DWT draws a random sample of size $n$ [43] from the two KDE distributions of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$, and selects a threshold $t$ that minimizes FPR and FNR. If a $t$ is found such that the observed FPR and FNR is 00 across multiple random sample draws, a reasonable statistical assurance can be given that the true FPR and FNR are both less than $n^{-1}$. Formally, if the observed FPR and FNR is 0 for $m=\lceil-\ln(1-\gamma)\rceil$ sample draws of $n$, a $\gamma$ confidence interval of $[0,\frac{1}{n})$ is achieved for the true FPR and FNR, where $\gamma\in[0,1)$ is the confidence level. We give a proof of this result, along with the theoretical analysis and explanation for each step of DWT in Appendix B.
By using KDE, we make minimal assumptions regarding the distribution of data. Therefore, DWT would work irrespective of whether the underlying data distribution is normal or not. To summarize, DWT encapsulates the following characteristics: (1) it requires as few as $4$ independent $\mathcal{M}_{clean}$ and $\mathcal{M}_{wm}$ to set a threshold (cf. Appendix B); (2) it gives statistical assurance of FPR and FNR being less than $n^{-1}$; (3) it generalizes to all data distributions and watermarking scheme using accuracy or $AUC_{\mathcal{D}}^{\mathcal{M}}$ as the basis for deciding ownership; and (4) it is robust to outliers encountered in the data, i.e., the threshold changes dynamically in response to outliers present in the data.
In a practical setting, $t$ only needs to be calculated by $\mathcal{J}$ in case of a dispute (cf. §4.4). Considering most models will not encounter ownership disputes, we consider the expected computational cost involved in the calculation of $t$ to be reasonable. In our experiments, we use 10 independent $\mathcal{M}_{clean}$ and $\mathcal{M}_{wm}$ instead of 4 to get tighter bounds on $t$. Figure 4 illustrates DWT using Yeast with SEAL for setting threshold. The threshold for each dataset-architecture pair using DWT is listed in TABLE III. Please note that a watermark threshold for a dataset-architecture pair appearing “low” (e.g., arXiv-GCN, NS-SEAL) may raise concern of OD falsely classifying $\mathcal{M}_{clean}$ as $\mathcal{M}_{wm}$, i.e., higher FPR. Despite the concern, the watermark thresholds can still be used for OD as DWT ensures FPR and FNR to be low. The verity of these “low” watermark thresholds can also be empirically justified from the corresponding $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ being “low” in TABLEs C.1, C.2, C.3, and C.4 in Appendix C.
Figure 4: KDEs of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ for dataset-architecture pair Yeast-SEAL, along with the corresponding watermark threshold $t$ with $n=10^{6}$. $\alpha$ and $\beta$ here denote the FPR and FNR (not visible since they are 00), while $\mu_{c}$ and $\mu_{w}$ denote the population mean of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$.